OpenStudy (anonymous):

how to find values for x and y in x+y=14 but x-y=2

OpenStudy (kl0723):

if it is a system then: x+y=14 x-y=2... positive y cancel with negative y... x=16... plug in x value in first eq and get (16)+y=14... y=14+16=30.

OpenStudy (anonymous):

Thanks ^^^